Total profit P is the difference between total revenue R and total cost C. Given the following​ total-revenue and​ total-cost functions, find the total​ profit, the maximum value of the total​ profit, and the value of x at which it occurs..
R(x)= 1200x-(x squared)
C(x)= 3100+20x

Answer by solve_for_x(190)

The profit function is:

P(x) = R(x) - C(x)

P(x) = (1200x - x^2) - (3100 + 20x)

P(x) = -x^2 + 1200x - 20x - 3100

P(x) = -x^2 + 1180x - 3100

This is the equation of a parabola that opens downward. The maximum value of the profit
corresponds to the vertex of the parabola.

The x-coordinate of the vertex is:

x = -b/2a = (-1180) / (-1*2) = 590

The function value at the vertex is then:

P(590) = -(590)^2 + 1198(590) - 3100

P(590) = 345,000

Solution:

Total profit = -x^2 + 1180x - 3100

Maximum profit = 345,000, at x = 590